A monopoly produces widgets at a marginal cost of $8 per unit and zero fixed costs. It faces an inverse demand function given by P = 38 - Q.
What are the profits of the monopoly in equilibrium?
A. $225
B. $120
C. $345
D. None of the statements associated with this question are correct
Answer #1
Is equilibrium defined as the maximum profit condition? You can have "equilbrium" at any price.

Is Q the quantity sold and P the price?

If so, profit is
Y = QP - 8Q = Q(P - 8) = Q (30 - Q)

maximum profit is abtained when dY/dQ = 0
30 = 2Q
Q = 15
Profit = 15 x 15 = $225
answered by: hello kitty